After a promotion, Gwen's annual salary was increased by 7.8% to R450000. Determine his monthly salary and the monthly increas he will be getting

1
Expert's answer
2022-02-17T12:24:30-0500

His monthly salary will be:

R450000/12 = R37500.

The old monthly salary was:

R37500/1.078 = R34786.64

The monthly increase he will be getting is:

37500 - 34786.64 = R2713.36.